devon purchased tickets to an air show for 7 adults and 2 children...the total cost was 92.00 the cost of the childs ticket was 8.00 less than the cost of an adults ticket...find the price of an adults ticket and a childs ticket?
Answer by checkley77(12844)

7(x+8)+2x=92
7x+56+2x=92
9x=92-56
9x=36
x=36/9
x=$4 ans. for the price of the children tickets.
4+8=$12 ans. for the price of the adult tickets.
Proof:
7*12+2*4=92
84+8=92
92=92